what is the formula to calculate the no. of onto functions from A to  B

Grade:12

8 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear prakyath

If A and B are two sets having m and n elements respectively such that   1≤n≤m  then number of onto function from A to B is

  ∑ (-1)n-r  nCr rm   r vary from 1 to n




Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Rehan Vijay Jain
11 years ago

hey can u tell me that in dis ques what shud i take the value of r

Q:- If A= {1,2,3,4}  B={4,5,6}, then find he no. of onto functions in this case.

pls tell me as fast it is possible

i need the answer of it too..........

Raawi Teza
11 years ago

No of elements of A = 4

No of elements of B = 3

usint this formula  ∑ (-1)n-r  nCr rm   r vary from 1 to n

{[(-1)^(3-1)]*3C1*1^4} + {[(-1)^(3-2)]*3C2*2^4{ +{ [(-1)^(3-3)]*3C3*3^4}

= 1*3*1 + (-1)*3*16 + 1*1*81

= 3-48+81

=36

namita
10 years ago

HI

 

For 

set a={a,b,c} and B={m,n}

 

the number of onto functions by your formula is 6

 

where as when i try manually it comes 8

 

according to you what should be the anwer

Abishek
10 years ago
If we calculate manually for number of onto functions from A=3 elements to B=2 elements, it comes out to be 6 as from the formula. Can anyone explain the formula.
Manas Shukla
7 years ago
If there is set A containing a elements and set B containing b elements, formula to calculate no of onto functions is : {(b to power a)- b}
Torun
6 years ago
n^r-nc1(n-1)^r+nc2(n-2)^r-nc3(n-3)^r.....where n=no. of elements in B;r=no. of elements in A. For a function ,f(A to B).Use this one guys ,its perfectttt.
ankit singh
3 years ago
Therefore, each element of X has 'n' elements to be chosen from. Therefore, total number of functions will be n×n×n.. m times = nm. For example: X = {a, b, c} and Y = {4, 5}